OpenStudy (anonymous):

What is the correct solution to this problem? A number divided by 7 is 6 less than –8. What is the number? A. –98 B. –56 C. –14 D. –7

OpenStudy (anonymous):

@Kpinky

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@countonme123

OpenStudy (johnweldon1993):

"A number divided by 7 is 6 less than –8." Let the number = 'x' So x divided by 7 is 6 less than -8 \[\large \frac{x}{7} = - 8 - 6\] \[\large \frac{x}{7} = -14\] \[\large x = -14 \times 7\]

OpenStudy (anonymous):

so @johnweldon1993 the answer is -14

OpenStudy (johnweldon1993):

No it is -14 times 7...I just wanted you to do that math there lol

OpenStudy (anonymous):

so its -98 @johnweldon1993

OpenStudy (johnweldon1993):

Correct!